a compound contains 4.07% of hydrogen ;24.27% of carbon and 71.65% of chlorine.Its molar mass is 98.96g then find the empirical formula
 
 
 

Grade:11

1 Answers

Arun
6 years ago
2Cl4H2ClEmpirical Formula Mass =12+2+35 = 49n = 98.96/46 = 2(almost)Molecular formula = n x empirical formula = C2Number of Moles of H = 4.07/1 = 4 (almost)Number of Moles of C = 24.27/12 = 2(almost)Number of Moles of Cl = 71/35.5 = 2​(almost)RatioC:H:Cl = 2:4:2 = 1:2:1Empirical Formula = CH